OpenStudy (anonymous):

Estimate the radius of a circle with the circumference of 272.2 inches. Use 3.14

OpenStudy (welshfella):

c = 2*pi* r make r the subject

OpenStudy (welshfella):

then plug in your values

OpenStudy (welshfella):

to make r the subject divide both sides by 2pi

OpenStudy (gabylovesyou):

r = C/2π. r = 272.2/(2 * pi) http://www.wolframalpha.com/input/?i=x+%3D+272.2%2F%282+*+pi%29 43.322 is the radius.

OpenStudy (anonymous):

Yes gabby is right

OpenStudy (anonymous):

I just did the math lol

OpenStudy (anonymous):

ok thank you. :-)